# SplatFlow: Self-Supervised Dynamic Gaussian Splatting in Neural Motion Flow Field for Autonomous Driving

### v1

@@ Abstract
Most existing Dynamic Gaussian Splatting methods for complex dynamic urban scenarios rely on accurate object-level supervision from expensive manual labeling, limiting their scalability in real-world applications. In this paper, we introduce SplatFlow, a Self-Supervised Dynamic Gaussian Splatting within Neural Motion Flow Fields (NMFF) to learn 4D space-time representations without requiring tracked 3D bounding boxes, enabling accurate dynamic scene reconstruction and novel view RGB/depth/flow synthesis. SplatFlow designs a unified framework to seamlessly integrate time-dependent 4D Gaussian representation within NMFF, where NMFF is a set of implicit functions to model temporal motions of both LiDAR points and Gaussians as continuous motion flow fields. Leveraging NMFF, SplatFlow effectively decomposes static background and dynamic objects, representing them with 3D and 4D Gaussian primitives, respectively. NMFF also models the status correspondences of each 4D Gaussian across time, which aggregates temporal features to enhance cross-view consistency of dynamic components. SplatFlow further improves dynamic scene identification by distilling features from 2D foundational models into 4D space-time representation. Comprehensive evaluations conducted on the Waymo Open Dataset and KITTI Dataset validate SplatFlow’s state-of-the-art (SOTA) performance for both image reconstruction and novel view synthesis in dynamic urban scenarios.
@@ 1 Introduction
As autonomous driving systems increasingly shift toward end-to-end models, there is a growing need for scalable simulation environments without domain gaps where these systems can undergo closed-loop evaluation. A promising approach is real-world closed-loop evaluation, which demands controllable sensor inputs, thereby driving the development of advanced scene reconstruction techniques. In this context, Neural Radiance Fields (NeRFs) [9] and 3D Gaussian Splatting (3DGS) [6] have proven to be effective in creating high-quality 3D scene reconstructions with excellent visual and geometric accuracy. However, accurately and comprehensively reconstructing dynamic driving scenes remains a major challenge, given the complexity of real-world scenarios without dynamic object annotations.
@@
Recent techniques enhance NeRF-based models by using object detection and tracking to animate dynamic objects, enabling photo-realistic view generation of dynamic urban street environments. Approaches [10, 19, 7, 22, 16, 3] create a scene graph, where both dynamic objects and the static background are represented as nodes, reconstructed within their canonical frames. Self-supervised methods [17, 24] model dynamic driving scenes by combining static neural field and time-dependent neural field to handle the static background and moving foreground objects separately. The SUDS [17] uses optical flow to ease the strict requirement for object labeling, while EmerNerf [24] learns temporal attributes and features in a self-supervision manner to minimize reliance on optical flow. However, despite utilizing implicit representations, these NeRF style methods face efficiency challenges in both training and inference, which becomes a major bottleneck for large-scale scene reconstruction and rendering tasks.
@@
While NeRF has proven effective for driving scenes, 3DGS offers a promising alternative due to faster training and rendering speed with more explicit representation. However, the original 3DGS faces notable challenges in modeling dynamic urban environments due to its limited representation capabilities in the temporal dimension. To address this, DrivingGaussian [27] introduces composite dynamic Gaussian graphs to handle multiple moving objects and incremental static Gaussians for background representation. StreetGaussian [23] optimizes the tracked bounding boxes of dynamic Gaussians along with 4D spherical harmonics to capture changing vehicle appearances. Both methods, however, require accurate object-level supervision, such as 3D object boxes and trackers, to decompose static and dynamic elements, limiting their scalability in real-world applications. In response, PVG [1] proposes a Periodic Vibration Gaussian mechanism to model dynamic driving scenes without relying on manually labeled 3D bounding boxes. However, the scene decomposition in PVG replies on time-dependent Gaussian attributes optimized only through rendering losses, but ignores motion information within the input point cloud for Gaussian initialization, leading to suboptimal performance in challenging scenarios with rapid movements.
To facilitate accurate scene reconstruction and real-time rendering without expensive annotations, we propose SplatFlow, a self-supervised Dynamic Gaussian Splatting method in Neural Motion Flow Fields (NMFF) for dynamic urban scenarios. In contrast to existing methods, as illustrated in Fig. 1, SplatFlow decomposes dynamic objects and the static background in a self-supervised way without requiring expensive 3D bounding box annotations. The key idea of SplatFlow is to seamlessly integrate 4D Gaussian representations within NMFF in a unified framework. The NMFF is a set of implicit functions to model temporal motions of both LiDAR points and Gaussians as continuous motion flow fields. Levaraging NMFF, SplatFlow not only enables the decomposition of dynamic and static elements from 3D LiDAR points, but also enables the status conversion of each 4D Gaussian across time. During Gaussian splatting, we represent dynamic objects using aggregated 4D Gaussians at various viewpoints and timestamps, and the static background using 3D Gaussians, respectively. In SplatFlow, NMFF’s capabilities are enhanced through three components: 1) learning 3D priors by pretraining on 3D LiDAR data, 2) optimizing temporal status transitions of 4D Gaussians on image data, and 3) distilling knowledge from 2D foundation models.
@@
  • NMFF models status correspondences of each 4D Gaussian over time, aggregating temporal features to enhance the cross-view consistency of dynamic components.
  • SplatFlow enhances the dynamic scene identification by uplifting 2D features distilled from foundational models to 4D space-time through optical flow rendering.
Comprehensive experiments on the Waymo [15] and KITTI [4] benchmarks demonstrate that SplatFlow outperforms the state-of-the-art (SOTA) methods in both image reconstruction and novel view synthesis for dynamic urban scenes. Notably, SplatFlow achieves this by avoiding tracked 3D bounding boxes of dynamic objects, allowing the proposed model to learn from extensive, in-the-wild data sources.
@@ 2 Related work
Real-world simulation is essential for generating data to support end-to-end autonomous driving solutions. However, current simulation engines like AirSim [14] and CARLA [2] encounter difficulties due to the expensive manual work needed to create virtual environments and the domain gap between real and simulated data. The rapid progress in Novel View Synthesis (NVS) technologies, such as NeRF [9] and 3DGS [6], enables 3D reconstruction and photorealistic image generation, significantly improving real-world simulation for autonomous driving.
@@
For 3D reconstruction and neural rendering in autonomous driving scenarios, sensor fusion solution combining surrounding cameras and LiDAR is commonly used in existing work [13, 8, 11, 5, 25]. Urban Radiance Field [13] improves NeRF training by incorporating 3D information from LiDAR to enhance 3D geometry learning. DNMP [8] employs a pre-trained deformable mesh primitive to represent the 3D scene, enhancing neural rendering quality. NPLF [11] uses explicit 3D reconstructions from LiDAR data to efficiently model the radiance field during rendering. StreetSurf [5] categorizes scenes into close-range, distant-view, and sky sections, achieving superior reconstruction of urban street surfaces. TCLC-GS [25] design a hybrid explicit and implicit 3D representation derived from LiDAR-camera data, to enrich the properties of 3D Gaussians for splatting. While these methods deliver impressive results for static urban scene rendering, they struggle with handling dynamic objects common in driving scenarios.
To model dynamic urban scenes, recent approaches [7, 10, 19, 17, 24] decompose the entire scene into static background and dynamic objects, learning their representations separately. PNF [7] uses monocular 3D bounding box predictions to isolate dynamic objects and further jointly optimizes their poses during the reconstruction process. NSG [10] uses neural graphs to represent entire scenes, decomposing dynamic multi-object environments. MARS [19] employs distinct sub-networks to model backgrounds and dynamic objects, creating an instance-aware simulation framework. These methods typically rely on manually annotated or predicted 3D bounding boxes. In order to avoid the need of 3D bounding boxes, SUDS [17] introduces a scalable hash table to represent large-scale dynamic urban scenes, using an off-the-shelf 2D optical flow estimator to track dynamic objects. EmerNerf [24] addresses this challenge by learning scene flow and using it to link corresponding points in the 4D neural field across multi-frame reconstruction, allowing the separation of static and dynamic objects without 3D bounding boxes. However, these methods, which rely on implicit representations, still suffer from inefficiency in both construction and rendering.
@@
Recently, 3DGS [6] introduced a novel explicit 3D scene representation, combining high-quality volume rendering with a fast speed. However, 3DGS is designed for static scenes and fails when modeling dynamic moving objects. To address this, DrivingGaussian [27] and StreetGaussian [23] partition the driving scene into static and dynamic components by separate sets of Gaussians based on vehicles’ 3D bounding boxes. DrivingGaussian [27] uses a composite dynamic Gaussian graph to manage multiple moving objects while incrementally employing incremental static 3D Gaussians to represent the entire background. StreetGaussian [23] models dynamic urban scenes as point clouds with semantic logits and 3D Gaussians, handling foreground vehicles and background separately. HUGS [26] also adopts 3D object boxes to identify dynamic elements from background, optimizing geometry, appearance, semantics, and motion specifically of dynamic Gaussians. However, all these methods rely on expensive annotated or predicted 3D bounding boxes to learn the time-dependent representations of dynamic objects. Most recent method PVG [1] introduces periodic vibration based Gaussian attributes, optimized through self-supervision without the need of 3D bounding boxes. Each Gaussian models dynamic changes over time through optimizable attributes including vibration directions, life span, and life peak.
@@ 3.1 Overview
By collecting images and point clouds with timestamps from surrounding cameras and LiDAR, we aim to learn a space-time 4D Gaussian representation of a dynamic scene without any human annotations, enabling fast and high-quality novel view rendering. As the pipeline of SplatFlow shown in Fig. 2, NMFF models the temporal motions of both 3D points and Gaussians as continuous motion flow fields. It serves two key roles in SplatFlow: 1) decomposing static and dynamic elements from 3D LiDAR points; 2) modeling the temporal status correspondences of each 4D Gaussian across time. We first decompose the time-series LiDAR points into static and dynamic points by NMFF, which are then separately merged to initialize the static and dynamic Gaussians. Dynamic objects are represented by aggregated 4D Gaussians associated with NMFF over time, while the static background is modeled using 3D Gaussians. NMFF learns the status correspondence of each 4D Gaussian in conjunction with optimizing 4D Gaussians’ attributes along time. We also distill optical flow knowledge from a 2D foundational model into 4D space-time representation. Finally, images, depths, and optical flows are rendered from novel viewpoints at different timesteps in the dynamic driving scenario.
@@ 3.2 Problem Definition:
$\Sigma^{\prime}(t)=JE\Sigma(t)E^{T}J^{T},\vspace{-2mm}$ (2)
where $E$ refers to the world-to-camera matrix and $J$ refers to Jacobian of the perspective transformation.
By sorting the Gaussians according to their depth within the camera space, we use $\alpha$-blending to estimate the color $C$ and depth $D$ at each pixel $p$ as,
$C=\sum_{i=1}^{N}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}),D=\sum_{i=1}^{N}z_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}).\vspace{-3mm}$ (3)
@@ 3.4 Neural Motion Flow Field
We first pre-train $\Phi$ on a sequence of 3D LiDAR points by forward and backward 3D geometry consistentency, enabling NMFF to learn the 3D motion of each query 3D point over time. Given the source point clouds $\mathcal{P}_{t_{1}}$ at time $t_{1}$ and target point clouds $\mathcal{P}_{t_{2}}$ at time $t_{2}$, we minimize the point distance between the source and target point clouds by a bidirectional Chamfer Distance (CD),
$CD(\mathcal{P}_{t_{1}},\mathcal{P}_{t_{2}})=\sum_{\mathclap{p_{t_{2}}\in\mathcal{P}_{t_{2}}}}D(p_{t_{2}},\mathcal{P}_{t_{1}})+\sum_{\mathclap{p_{t_{1}}\in\mathcal{P}_{t_{1}}}}D(p_{t_{1}},\mathcal{P}_{t_{2}}),\vspace{-2mm}$ (5)
where $D$ refers to point distance function, where correspondences from source-to-target and target-to-source are searched among the nearest point neighbors.
We generate a 3D dynamic mask to separate static and dynamic points through ego-motion compensation on the scene flow predicted by NMFF. Static and dynamic point clouds are then merged separately using ego-motion and scene flow. For the static background, we initialize the static Gaussians using the merged 3D static points. For dynamic objects, we aggregate points from different timestamps into a common reference frame using scene flow predictions. This aggregation densifies and completes dynamic point clouds from various viewpoints over time.
4D Gaussian within NMFF: The pre-trained $\Phi$ is subsequently integrated with 4D Gaussian representations of dynamic objects, enabling joint optimization during Gaussian splatting. Given consecutive time $t_{1}$ and $t_{2}$ where $t_{1}<t_{2}$ with respective states $\{\mathcal{G}_{i}(t_{1})\}$ and $\{\mathcal{G}_{i}(t_{2})\}$, these states are connected by NMFF for each Gaussian as,
$\mathcal{G}_{i}(t_{1})=\{\mu(t_{1}),R(t_{1}),S,\alpha,c\},\vspace{-1mm}$ (6)
$\Delta\mu_{t_{1}:t_{2}},~{}\Delta R_{t_{1}:t_{2}}=\phi_{t_{1}:t_{2}}(\mu(t_{1})),\vspace{-0.5mm}$ (7)
$\hat{\mathcal{G}_{i}}(t_{2})=\{\mu(t_{1})+\Delta\mu_{t_{1}:t_{2}},R(t_{1})\cdot\Delta R_{t_{1}:t_{2}},S,\alpha,c\}.$ (8)
@@
We employ NMFF $\Phi$ correspondence to warp the center of all 4D Gaussians across time to the desired timestamp, as $\{\mathcal{G}_{i}(t)\}=\{\mathcal{G}(t),\hat{\mathcal{G}}(t_{1}),\hat{\mathcal{G}}(t_{2}),...,\hat{\mathcal{G}}(t_{n})\}$. For the transition of Gaussians across multiple timestamps, we apply the corresponding NMFF through step-by-step propogation. Here NMFF learns a motion transition of the 4D Gaussian among training frames in a self-supervised manner, promoting a multi-view consistent representation.
For static background, we represent it by static 3D Gaussians $\mathcal{G}_{static}$ following the strategy [1]. For sky area, we adopt a separate environmental map with sky Gaussians $\mathcal{G}_{sky}$ following [1, 24]. Given a specific viewpoint $[E_{t},I_{t},t]$, the image $\hat{\mathcal{I}}$ and depth $\hat{\mathcal{D}}$ at time $t$ are rendered by the differentiable rendering using a set of Gaussians,

### v2

@@ Abstract
Most existing Dynamic Gaussian Splatting methods for complex dynamic urban scenarios rely on accurate object-level supervision from expensive manual labeling, limiting their scalability in real-world applications. In this paper, we introduce SplatFlow, a Self-Supervised Dynamic Gaussian Splatting within Neural Motion Flow Fields (NMFF) to learn 4D space-time representations without requiring tracked 3D bounding boxes, enabling accurate dynamic scene reconstruction and novel view RGB/depth/flow synthesis. SplatFlow designs a unified framework to seamlessly integrate time-dependent 4D Gaussian representation within NMFF, where NMFF is a set of implicit functions to model temporal motions of both LiDAR points and Gaussians as continuous motion flow fields. Leveraging NMFF, SplatFlow effectively decomposes static background and dynamic objects, representing them with 3D and 4D Gaussian primitives, respectively. NMFF also models the correspondences of each 4D Gaussian across time, which aggregates temporal features to enhance cross-view consistency of dynamic components. SplatFlow further improves dynamic object identification by distilling features from 2D foundation models into 4D space-time representation. Comprehensive evaluations conducted on the Waymo and KITTI Datasets validate SplatFlow’s state-of-the-art (SOTA) performance for both image reconstruction and novel view synthesis in dynamic urban scenarios.
@@ 1 Introduction
As autonomous driving systems increasingly shift toward end-to-end models, there is a growing need for scalable simulation environments without domain gaps where these systems can undergo closed-loop evaluations. A promising approach is real-world closed-loop evaluation, which demands controllable sensor inputs, thereby driving the development of advanced scene reconstruction techniques. In this context, Neural Radiance Fields (NeRFs) [10] and 3D Gaussian Splatting (3DGS) [7] have proven to be effective in creating high-quality 3D scene reconstructions with excellent visual and geometric accuracy. However, accurately and comprehensively reconstructing dynamic driving scenes remains a major challenge, given the complexity of real-world scenarios without dynamic object annotations.
@@
Recent approaches advance NeRF-based representations by incorporating object detection and tracking to model dynamic elements, enabling photo-realistic view generation of dynamic urban street environments. Approaches [12, 22, 8, 24, 19, 3] create a scene graph, where both dynamic objects and the static background are represented as nodes, reconstructed within their canonical frames. Self-supervised methods [20, 26] model dynamic driving scenes by combining static neural field and time-dependent neural field to handle the static background and moving foreground objects separately. The SUDS [20] uses optical flow to ease the strict requirement for object labeling, while EmerNerf [26] learns temporal attributes and features in a self-supervision manner to avoid reliance on optical flow. However, these NeRF style methods face efficiency challenges in both training and inference, which becomes a major bottleneck for large-scale scene reconstruction and rendering.
@@
While NeRF has proven effective for driving scenes, 3DGS offers a promising alternative due to faster training and rendering speed with more explicit representations. However, the original 3DGS faces notable challenges in modeling dynamic urban environments due to its limited representation capabilities in the temporal dimension. To address this, DrivingGaussian [30] introduces composite dynamic Gaussian graphs to handle multiple moving objects and incremental static Gaussians for background representation. StreetGaussian [25] optimizes the tracked bounding boxes of dynamic Gaussians along with 4D spherical harmonics to capture changing vehicle appearances. Both methods, however, require accurate object-level supervision, such as 3D object boxes and trackers, to decompose static and dynamic elements, limiting their scalability in real-world applications. In response, PVG [1] proposes a Periodic Vibration Gaussian mechanism to model dynamic driving scenes without relying on manually labeled 3D bounding boxes. However, the scene decomposition in PVG relies on time-dependent Gaussian attributes optimized by rendering losses, but ignores motion cues within the input point cloud used for Gaussian initialization, leading to suboptimal performance in challenging scenarios with rapid movements. To facilitate accurate scene reconstruction and real-time rendering without expensive annotations, we propose SplatFlow, a self-supervised Dynamic Gaussian Splatting method in Neural Motion Flow Fields (NMFF) for dynamic urban scenarios. In contrast to existing methods, as illustrated in Fig. 1, SplatFlow decomposes dynamic objects and the static background in a self-supervised way without requiring expensive 3D bounding box annotations. The key idea of SplatFlow is to seamlessly integrate 4D Gaussian representations within NMFF in a unified framework. The NMFF is a set of implicit functions to model temporal motions of both LiDAR points and Gaussians as continuous motion flow fields. Levaraging NMFF, SplatFlow not only enables the decomposition of dynamic and static elements from 3D LiDAR points, but also enables the status conversion of each 4D Gaussian across time. During Gaussian splatting, we represent dynamic objects using aggregated 4D Gaussians at various viewpoints and timestamps, and the static background using 3D Gaussians, respectively. In SplatFlow, NMFF’s capabilities are enhanced through three components: 1) learning 3D motion priors by pretraining on 3D LiDAR data, 2) optimizing temporal status transitions of 4D Gaussians on image data, and 3) distilling knowledge from 2D foundation models.
@@
  • NMFF models correspondences of each 4D Gaussian over time, aggregating temporal features to enhance the cross-view consistency of dynamic components.
  • SplatFlow enhances the dynamic object identification by uplifting 2D features distilled from foundation models to 4D space-time through optical flow rendering.
Comprehensive experiments on the Waymo [17] and KITTI [4] benchmarks demonstrate that SplatFlow outperforms the state-of-the-art (SOTA) methods in both image reconstruction and novel view synthesis for dynamic urban scenes. Notably, SplatFlow achieves this without relying on tracked 3D bounding boxes of dynamic objects, allowing the proposed model to learn from extensive, in-the-wild data sources.
@@ 2 Related work
Real-world simulation is essential for generating data to support end-to-end autonomous driving solutions. However, current simulation engines like AirSim [16] and CARLA [2] encounter difficulties due to the expensive manual work needed to create virtual environments and the domain gap between real and simulated data. The rapid progress in Novel View Synthesis (NVS) technologies, such as NeRF [10] and 3DGS [7], and following work  [11, 27, 18, 6] enables 3D reconstruction and photorealistic image generation, significantly improving real-world simulation for autonomous driving. For 3D reconstruction and neural rendering in autonomous driving scenes, sensor fusion solution combining surrounding cameras and LiDAR is commonly used in existing work [15, 9, 13, 5, 28]. Urban Radiance Field [15] improves NeRF training by incorporating 3D data from LiDAR to enhance 3D geometry learning. DNMP [9] employs a pre-trained deformable mesh primitive to represent the 3D scene, enhancing neural rendering quality. NPLF [13] uses explicit 3D reconstructions from LiDAR data to efficiently model the radiance field. TCLC-GS [28] design a hybrid explicit and implicit 3D representation derived from LiDAR-camera data, to enrich the properties of 3D Gaussians for splatting. Though these methods show impressive results for static urban scene rendering, they struggle with dynamic objects common in driving scenarios.
@@
To model dynamic urban scenes, recent approaches [8, 12, 22, 20, 26] decompose the entire scene into static background and dynamic objects, learning their representations separately. PNF [8] uses monocular 3D bounding box predictions to isolate dynamic objects and further jointly optimizes their poses during the reconstruction process. NSG [12] uses neural graphs to represent entire scenes, decomposing dynamic multi-object environments. MARS [22] employs distinct sub-networks to model backgrounds and dynamic objects, creating an instance-aware simulation framework. These methods typically rely on manually annotated or predicted 3D bounding boxes. In order to avoid the need of 3D bounding boxes, SUDS [20] introduces a scalable hash table to represent large-scale dynamic urban scenes, using an off-the-shelf 2D optical flow estimator to track dynamic objects. EmerNerf [26] addresses this challenge by learning scene flow to associate points across time in the 4D neural radiance field, allowing the separation of static and dynamic objects without 3D bounding boxes. However, these methods, which rely on implicit representations, still suffer from inefficiency in both reconstruction and rendering. Recently, 3DGS [7] introduced a novel explicit 3D scene representation, combining high-quality volume rendering with a fast speed. However, 3DGS is designed for static scenes and fails when modeling dynamic moving objects. To address this, DrivingGaussian [30] and StreetGaussian [25] partition the driving scene into static and dynamic components by separate sets of Gaussians based on the 3D bounding boxes of vehicles. DrivingGaussian [30] uses a composite dynamic Gaussian graph to manage multiple moving objects while incrementally employing incremental static 3D Gaussians to represent the static background. StreetGaussian [25] models foreground vehicles with optimizable poses and a 4D spherical harmonics appearance model, and background with 3D Gaussians separately. HUGS [29] also adopts 3D object boxes to identify dynamic elements, enabling optimization of geometry, appearance, semantics, and motion attributes for dynamic Gaussians. However, all these methods rely on expensive annotated or predicted 3D bounding boxes to learn the time-dependent representations of dynamic objects. Most recent method PVG [1] introduces periodic vibration based Gaussian attributes, optimized through self-supervision without the need of 3D bounding boxes. Each Gaussian models dynamic changes over time through optimizable attributes including vibration directions, life span, and life peak.
@@ 3.1 Overview
By collecting images and point clouds with timestamps from surrounding cameras and LiDAR, we aim to learn a space-time 4D Gaussian representation of a dynamic scene without any human annotations, enabling fast and high-quality novel view rendering. As the pipeline of SplatFlow shown in Fig. 2, NMFF models the temporal motions of both 3D points and Gaussians as continuous motion flow fields. It serves two key roles in SplatFlow: 1) decomposing static and dynamic elements from 3D LiDAR points; 2) modeling the temporal correspondences of each 4D Gaussian across time. We first decompose the time-series LiDAR points into static and dynamic points by NMFF, which are then separately merged to initialize the static and dynamic Gaussians. Dynamic objects are represented by aggregated 4D Gaussians associated with NMFF over time, while the static background is modeled using 3D Gaussians. NMFF learns the correspondence of each 4D Gaussian in conjunction with optimizing 4D Gaussians’ attributes along time. We also distill optical flow knowledge from a 2D foundation model into 4D space-time representation. Finally, images, depths and optical flows are rendered from novel viewpoints at different timesteps in dynamic driving scenes.
@@ 3.2 Problem Definition:
$\Sigma^{\prime}(t)=JE\Sigma(t)E^{T}J^{T},\vspace{-1mm}$ (2)
where $E$ refers to the world-to-camera matrix and $J$ refers to Jacobian of the perspective transformation. By sorting the Gaussians according to their depth within the camera space, we use $\alpha$-blending to estimate the color $C$ and depth $D$ at each pixel $p$ as,
$C=\sum_{i=1}^{N}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}),D=\sum_{i=1}^{N}z_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}).\vspace{-2mm}$ (3)
@@ 3.4 Neural Motion Flow Field
We first pre-train $\Phi$ on a sequence of 3D LiDAR points by forward and backward 3D geometry consistency, enabling NMFF to learn the 3D motion of each query 3D point over time. Given the source point clouds $\mathcal{P}$ and target point clouds $\mathcal{Q}$ from different timestamps, we minimize the point distance between the source and target point clouds by a bidirectional Chamfer Distance ($CD$),
$d_{\mathrm{CD}}(\mathcal{P},\mathcal{Q})=\sum_{p\in\mathcal{P}}\min_{q\in\mathcal{Q}}\|p-q\|^{2}+\sum_{q\in\mathcal{Q}}\min_{p\in\mathcal{P}}\|q-p\|^{2}.\vspace{-2mm}$ (5)
The bidirectional $CD$ is a point-based distance that establishes correspondences in both directions—source to target and target to source—by searching for the nearest neighbors in each point cloud.
We compensate for ego-car motion in LiDAR points to let NMFF learn only 3D motion flow during pre-training. As a result, the 3D motion flow predicted by NMFF excludes the ego-car motion. We generate a 3D dynamic mask to separate static and dynamic points by applying a threshold on each point’s 3D motion flow. We then merge static point clouds using ego-motion alone, and dynamic point clouds using both ego-motion and motion flow. For the static background, we initialize the static 3D Gaussians using the merged 3D static points. For dynamic objects, we initialize 4D Gaussians by aggregating points from different timestamps into a common reference frame (the mid-timestamp reference frame), which increases density and completeness by integrating observations from various viewpoints over time. 4D Gaussian within NMFF: The pre-trained $\Phi$ is subsequently integrated with 4D Gaussian representations of dynamic objects, enabling joint optimization during Gaussian splatting. Given consecutive time $t_{1}$ and $t_{2}$ where $t_{1}<t_{2}$ with respective states $\{\mathcal{G}_{i}(t_{1})\}$ and $\{\mathcal{G}_{i}(t_{2})\}$, these states are connected by NMFF for each Gaussian as,
$\mathcal{G}_{i}(t_{1})=\{\mu(t_{1}),R(t_{1}),S,\alpha,c\},\vspace{-3mm}$ (6)
$\Delta\mu_{t_{1}:t_{2}},~{}\Delta R_{t_{1}:t_{2}}=\phi_{t_{1}:t_{2}}(\mu(t_{1})),\vspace{-3mm}$ (7)
$\hat{\mathcal{G}_{i}}(t_{2})=\{\mu(t_{1})+\Delta\mu_{t_{1}:t_{2}},R(t_{1})\cdot\Delta R_{t_{1}:t_{2}},S,\alpha,c\}.\vspace{-2mm}$ (8)
@@
We employ NMFF $\Phi$ correspondence to warp the center of all aggregated 4D Gaussians across time to the desired timestamp, as $\{\mathcal{G}_{i}(t)\}=\{\mathcal{G}(t),\hat{\mathcal{G}}(t_{1}),\hat{\mathcal{G}}(t_{2}),...,\hat{\mathcal{G}}(t_{n})\}$. For the transition of Gaussians across multiple timestamps, we apply the corresponding NMFF through step-by-step propogation. In practice, we propogate the aggregated 4D Gaussian at the mid-timestamp to the target timestamp. Here NMFF learns a motion transition of the 4D Gaussian among training frames in a self-supervised manner, promoting a multi-view consistent representation. For static background, we represent it by static 3D Gaussians $\mathcal{G}_{static}$ following the strategy [1]. For sky area, we adopt a separate environmental map with static sky Gaussians $\mathcal{G}_{sky}$ following [1, 26]. Given a specific viewpoint $[E_{t},I_{t},t]$, the image $\hat{\mathcal{I}}$ and depth $\hat{\mathcal{D}}$ at time $t$ are obtained by the differentiable rendering using a set of Gaussians,